# Modality Inflation: Energy Characterization and Optimization Opportunities for MLLM Inference

### v1

@@ 3.5. Input Complexity
Figure 7. Impact of image resolution on energy, latency, and visual token generation in multimodal LLM inference.
@@
4. Discussion: Stage-wise GPU Frequency Scaling
Our experiments demonstrate that the energy consumption and efficiency of MLLM inference vary widely based on the characteristics of the inputs, the vision encoder structure, and the extent of visual token expansion. This suggests that naive, one-size-fits-all serving policies for MLLM inference lead to inefficient use of hardware resources, highlighting the need for energy-efficient and MLLM-specific serving mechanisms. We select InternVL3 and Qwen2.5-VL as case studies because they represent two contrasting energy bottlenecks in multimodal inference. Qwen2.5-VL is dominated by a compute-intensive vision encoder, whereas InternVL3 exhibits lower encoding cost but substantial prefill energy overhead. Hence, in this discussion, our main focus would be on the vision encoding and prefill energy components that dominate the energy overhead of multimodal inputs. We do not consider the energy of the decoding components, as the related energy has been thoroughly studied in the text-based LLM with scalability to the length of the outputs (liu2025greenllm). From both models, the energy per request heatmaps in each stage in Figure 8 demonstrate that a single global serving method where the GPU runs at its highest frequency does not necessarily result in minimum energy consumption per request during multimodal inference tasks. In these experiments, the energy heatmaps show that in most stages, the energy per request has a tendency to be at its minimum at intermediate frequencies rather than at the highest clock rate, which distinguishes latency optimal points from energy optimal points. Starting with InternVL3, the encode-stage heatmaps clearly indicate that significant opportunities for energy savings exist with limited performance impact. At a batch size of 32, increasing the GPU frequency from 1050 MHz to 1410 MHz reduces encoding latency from 0.18 s to 0.16 s (an 11.8% reduction), while throughput improves by 13.4%; however, energy per request increases from 1.03 J to 1.28 J (a 24.9% increase). The encode energy heatmaps further show that energy remains close to its minimum at intermediate frequencies, even as latency continues to decrease at higher clocks. This behavior indicates diminishing latency returns for encoding at higher frequencies and suggests that adjusting the encoding frequency offers meaningful opportunities for energy savings with modest performance loss, particularly when downstream stages dominate end-to-end latency.
@@ 4. Discussion: Stage-wise GPU Frequency Scaling
Figure 8. DVFS frequency–batch size heatmaps characterizing per-request energy behavior in the encoding and prefill stages of multimodal LLM inference.
5. Related Work
5.1. Energy Optimization for LLMs
Existing energy optimization techniques for large language models focus on text-only workloads and primarily propose optimizations for the prefill and decode stages of inference. These works explore a range of techniques, such as scheduling, batching, and power management, to reduce inference cost while maintaining acceptable latency (liu2025greenllm; 10946802; fernandez2025energy; wilkins2024offline). However, these approaches do not consider multimodal models and requests, where additional input modalities and encoder can significantly alter inference-stage behavior and introduce new sources of energy inefficiency. Unlike prior energy characterization studies that focus on text-only LLMs or earlier multimodal models under different hardware and workload assumptions, our work emphasizes controlled, within-study comparisons across modern vision–language models.
5.2. Efficient Serving Systems for MLLMs
Current serving methodologies for MLLMs are mainly focused on improved throughput and lower end-to-end latency using techniques such as efficient batching, pipelined parallelism, or hardware/model disaggregation (qiumodserve; liu2025elasticmm; dong2025hydrainfer; singh2024efficiently; guo2025rserve). While these methodologies have achieved great performance improvements, they do not focus on energy efficiency. As a result, existing methodologies have not attempted to analyze energy tradeoffs among different elements in multimodal inference, including a vision encoder and a language model backbone, nor have they explored energy-efficient hardware control for multimodal inference.
5.3. Power Management and DVFS in GPU-Based Systems
Power management in both GPU-based systems and datacenter settings has been studied widely, with a focus on improving power efficiency under performance requirements. In existing work, a variety of approaches have been considered for managing power usage, such as DVFS, power capping, and hardware transformations, which have common usage in dealing with latency or throughput requirements (patel2024characterizing; maliakel2025investigating). The above approaches generally rely on characteristics of workloads, in addition to system state information, for making decisions on power management in order to lower power consumption without overstepping performance requirements.
6. Future Work

### v2

@@ III-E Input Complexity
Figure 7: Impact of image resolution on energy, latency, and visual token generation in multimodal LLM inference.
@@
IV Discussion: Stage-wise GPU Frequency Scaling
Our experiments demonstrate that the energy consumption and efficiency of MLLM inference vary widely based on the characteristics of the inputs, the vision encoder structure, and the extent of visual token expansion. This suggests that naive, one-size-fits-all serving policies for MLLM inference lead to inefficient use of hardware resources, highlighting the need for energy-efficient and MLLM-specific serving mechanisms. We select InternVL3 and Qwen2.5-VL as case studies because they represent two contrasting energy bottlenecks in multimodal inference. Qwen2.5-VL is dominated by a compute-intensive vision encoder, whereas InternVL3 exhibits lower encoding cost but substantial prefill energy overhead. Hence, in this discussion, our main focus is on the vision encoding and prefill energy components that dominate the energy overhead of multimodal inputs. We do not consider the energy of the decoding components, as the related energy has been thoroughly studied in the text-based LLM with scalability to the length of the outputs [20]. From both models, the energy per request heatmaps in each stage in Figure 8 demonstrate that a single global serving method where the GPU runs at its highest frequency does not necessarily result in minimum energy consumption per request during multimodal inference tasks. In these experiments, the energy heatmaps show that in most stages, the energy per request has a tendency to be at its minimum at intermediate frequencies rather than at the highest clock rate, which distinguishes latency optimal points from energy optimal points. Starting with InternVL3, the encode-stage heatmaps clearly indicate that significant opportunities for energy savings exist with limited performance impact. At a batch size of 32, increasing the GPU frequency from 1050 MHz to 1410 MHz reduces encoding latency from 0.18 s to 0.16 s (an 11.8% reduction), while throughput improves by 13.4%; however, energy per request increases from 1.03 J to 1.28 J (a 24.9% increase). The encode energy heatmaps further show that energy remains close to its minimum at intermediate frequencies, even as latency continues to decrease at higher clocks. This behavior indicates diminishing latency returns for encoding at higher frequencies and suggests that adjusting the encoding frequency offers meaningful opportunities for energy savings with modest performance loss, particularly when downstream stages dominate end-to-end latency.
@@ IV Discussion: Stage-wise GPU Frequency Scaling
V Related Work
V-A Energy Optimization for LLMs
Existing energy optimization techniques for large language models focus on text-only workloads and primarily propose optimizations for the prefill and decode stages of inference. These works explore a range of techniques, such as scheduling, batching, and power management, to reduce inference cost while maintaining acceptable latency [20, 37, 9, 40]. However, these approaches do not consider multimodal models and requests, where additional input modalities and encoder can significantly alter inference-stage behavior and introduce new sources of energy inefficiency. Unlike prior energy characterization studies that focus on text-only LLMs or earlier multimodal models under different hardware and workload assumptions, our work emphasizes controlled, within-study comparisons across modern vision–language models.
V-B Efficient Serving Systems for MLLMs
Current serving methodologies for MLLMs are mainly focused on improved throughput and lower end-to-end latency using techniques such as efficient batching, pipelined parallelism, or hardware/model disaggregation [30, 21, 7, 35, 11]. While these methodologies have achieved great performance improvements, they do not focus on energy efficiency. As a result, existing methodologies have not attempted to analyze energy tradeoffs among different elements in multimodal inference, including a vision encoder and a language model backbone, nor have they explored energy-efficient hardware control for multimodal inference.
V-C Power Management and DVFS in GPU-Based Systems
Power management in both GPU-based systems and datacenter settings has been studied widely, with a focus on improving power efficiency under performance requirements. In existing work, a variety of approaches have been considered for managing power usage, such as DVFS, power capping, and hardware transformations, which have common usage in dealing with latency or throughput requirements [28, 23]. The above approaches generally rely on characteristics of workloads, in addition to system state information, for making decisions on power management in order to lower power consumption without overstepping performance requirements.
@@ V-C Power Management and DVFS in GPU-Based Systems
Figure 8: DVFS frequency–batch size heatmaps characterizing per-request energy behavior in the encoding and prefill stages of multimodal LLM inference.
VI Future Work
@@ VII Conclusion
Acknowledgment
This material is based on work supported by the National Science Foundation under Grants No 2315851 and 2106634 and the Virginia Tech Foundation.